mathslover (mathslover):

Logarithm Question (2) \(\log_{\sqrt{2} \sin x} (1-\sin x) = 2 \)

mathslover (mathslover):

@ParthKohli @myininaya @jim_thompson5910

OpenStudy (tkhunny):

What? Is that \(\sqrt{2}\sin(x)\) the Base of the Logarithm?

mathslover (mathslover):

Yes.

OpenStudy (anonymous):

ITs the base

mathslover (mathslover):

There are 2 methods that point towards a single result : \(\bf{\bullet}\) First Method : \(\log_{\sqrt{2} \sin x} (1-\sin x) = \log_{\sqrt{2} \sin x} ( \sqrt{2} \sin x)^2 \) Therefore, \(1-\sin x = 2\sin^2 x\)

OpenStudy (tkhunny):

Weird. Don't think I've ever seen that. You seem to have transformed it to the equivalent exponential form. That's good.

Parth (parthkohli):

Whoops, sorry.

Parth (parthkohli):

First, keep the conditions for a logarithm to be defined in your mind. Solve it only then.

mathslover (mathslover):

\(\bf{\bullet}\) Second Method : \(\cfrac{1}{2} \log_{\sqrt{2}\sin x} (1-\sin x) = 1 \\ \implies \log_{{\left(\sqrt{2} \sin x\right) }^2} (1-\sin x) = 1 \\ \implies \log_{2\sin^2 x } (1-\sin x) = 1 \\ \implies 1 -\sin x = 2 \sin^2 x \)

Parth (parthkohli):

Yes, solve the quadratic.\[\sqrt{2}\sin(x) > 0 \implies \sin(x) > 0,\sqrt{2}\sin(x) \ne 1\]\[1 - \sin(x) > 0 \implies \sin(x) < 1 \checkmark \]

OpenStudy (tkhunny):

Time!! I want to think about Domain issues. I always want to think about that. The definition of a base requires that it not be one (1). Your budy \(\sqrt{2}\sin(x)\) is often one (1). What's your plan for resolving that?

mathslover (mathslover):

We have an equation now : \(1-\sin x = 2\sin ^2 x\) Again, this can be solved in 2 ways : \(\bf{\bullet}\) First Method : \(\sin x = y\) ; (Say) So, \(1 - y = 2y^2 \\ \implies 2y^2 + y - 1=0 \) Using Quadratic Equation formula : \(y = \cfrac{-1 \pm \sqrt{9}}{4} = \cfrac{-1 \pm 3}{4} \) \(y = \cfrac{1}{2}\) or \(y = -1\)

Parth (parthkohli):

Why not just factor? :P\[\sin(x) > 0\]

mathslover (mathslover):

Well, if I take it directly @tkhunny : From log properties : 1 - sinx > 0 \(\implies x < \cfrac{\pi}{2}\) As, \(\sqrt{2} \sin x \ne 1\) Therefore, \(x \ne \cfrac{\pi}{4}\)

mathslover (mathslover):

And from the first method we got y = 1/2 or y = -1 That is \(\sin x = 1/2\) or \(\sin x = -1\)

Parth (parthkohli):

You don't need to do those. You have the values for \(\sin(x)\), so there is no need to find the angles for the conditions.

Parth (parthkohli):

Just eliminate \(\sin(x) = -1\) as \(\sin(x) >0\).

OpenStudy (tkhunny):

Where did we get \(\sin(x) > 0\)? That makes no sense. Did I miss an assumption along the way?

mathslover (mathslover):

@tkhunny , the base.

mathslover (mathslover):

\(\sqrt{2} \sin x > 0\)

mathslover (mathslover):

implies that : sin x > 0

mathslover (mathslover):

So, sin x can not be equal to 1. Thus, we have : sin x = 1/2 x = pi/6

mathslover (mathslover):

For the second method to solve 1 - sin x = \(2\sin^2 x\) We can do it like this also : \(1-2\sin^2 x = \sin x\) \(\cos (2x) = \sin x\)

mathslover (mathslover):

My main question, was just a very little equation (above) : cos(2x) = sin x How to conclude the answer from here?

OpenStudy (anonymous):

double angle formula

OpenStudy (anonymous):

cos(2x) = cos^2(x) - sin^2(x) and cos^2(x) = 1 - sin^2(x)

mathslover (mathslover):

I meant to say that, how to get the value of "x" from cos(2x) = sin(x)

OpenStudy (tkhunny):

I was arguing that with myself for a little while. I yield.

OpenStudy (anonymous):

everything now is in terms of sin(x). Solve it

OpenStudy (anonymous):

cos(2x) = sin(x) (1 - sin^2(x)) - sin^2(x) = sin(x)

mathslover (mathslover):

Yes, we can solve it by 1 - sin x = 2sin^2 x but, I wanted to know that is there any way that I can solve it directly from here : cos(2x) = sin (x) ?

mathslover (mathslover):

without simplifying it again into the original form (1-sin x = 2sin^2 x)

OpenStudy (anonymous):

well i don't know what you mean by solving directly. As far as I know, there is now way you can solve it without using trig identity. You could perhaps take the arccosine for both sides, but then you would need to know the properties of composition of trig inverse functions. This seems to make things rather more complicated

OpenStudy (anonymous):

there is *no* way

mathslover (mathslover):

Oh... thanks @sourwing ... I will prefer not to make it more complicated, and as I have not learnt trig. inverse functions, so, I will not try that either. :)

mathslover (mathslover):

Thank you @tkhunny @sourwing @ParthKohli .

OpenStudy (anonymous):

and there is actually infinitely many solutions to original equation ^^ https://www.wolframalpha.com/input/?i=log_%7Bsqrt%282%29+sin%28x%29%7D+%281+-+sin%28x%29%29+%3D+2

Parth (parthkohli):

You can try to see the unit circle.